two spherical objects have equal masses and experience a gravitational force of 25 N towards one other . Their centers are 36 cm apart . determine each of their masses .


1
Expert's answer
2022-03-25T15:02:20-0400

The Newton's law of universal gravitation says

"F=G\frac{m_1m_2}{r^2}=G\frac{m^2}{r^2}"

Hence, the mass of bodies

"m_1=m_2=\sqrt{\frac{Fr^2}{G}}"

"=\sqrt{\frac{25*0.36^2}{6.67*10^{-11}}}=2.2*10^5\:\rm kg"